IF 14000 is invested at 10% simple annual interest, how much should be invested at 12% annual simple interest so that the total yearly income from both investments is 5000?
Answer by Nate(3500)

I = PRT = 14000 * 0.1 * 1 = 1400
5000 - 1400 = 3600
so ....
3600 = P * 0.12 * 1
3600/0.12 = P
Your principle should be $30,000.00